# On Robust Reinforcement Learning with Lipschitz-Bounded Policy Networks

### v1

@@ Abstract
This paper presents a study of robust policy networks in deep reinforcement learning. We investigate the benefits of policy parameterizations that naturally satisfy constraints on their Lipschitz bound, analyzing their empirical performance and robustness on two representative problems: pendulum swing-up and Atari Pong. We illustrate that policy networks with small Lipschitz bounds are significantly more robust to disturbances, random noise, and targeted adversarial attacks than unconstrained policies composed of vanilla multi-layer perceptrons or convolutional neural networks. Moreover, we find that choosing a policy parameterization with a non-conservative Lipschitz bound and an expressive, nonlinear layer architecture gives the user much finer control over the performance-robustness trade-off than existing state-of-the-art methods based on spectral normalization.
@@
reinforcement learning, robustness, Lipschitz networks
@@ 1 Introduction
Deep reinforcement learning (Deep RL) has been the driving force behind many recent successes in learning-based control, including in discrete game-like problems [17], robotic manipulation [12], and locomotion [22]. However, the applicability of Deep RL to performance- and safety-critical systems is currently limited by questions of its robustness [11]. It is well known that neural networks can be highly sensitive to small input perturbations, making policy networks learned via Deep RL potentially unrobust to disturbances, noise, and targeted adversarial attacks [25, 6]. Despite sharing similar sensitivity issues to neural classifiers, for which many robust neural networks have recently been developed [3, 28, 30], the use of robust policy architectures in Deep RL has not been widely studied.
@@
The true $\ell_{2}$ Lipschitz constant, denoted $\mathrm{Lip}(f)$, is the smallest $\gamma$ satisfying (1). Lipschitz-bounded networks are then “smoother” and less sensitive to input perturbations since small variations to their inputs will induce only small variations in their outputs. Despite the wealth of recent work on constructing Lipschitz-bounded deep networks [3, 28, 30], to the best of our knowledge, the only method tested so far in Deep RL is spectral normalization [4, 26] which is known to give conservative bounds on the true Lipschitz constant of the network.
@@ 2.1 Deep reinforcement learning
$\max_{\theta}\ R:=\mathop{\mathbb{E}}\left[\sum_{t=0}^{\infty}\rho^{t}r_{t}(x_{t},u_{t})\right]\;\text{s.t.}\;\ x_{t+1}=f(x_{t},u_{t},w_{t}),\ u_{t}=\kappa(x_{t};\theta),$ (3)
@@ 2.2 Adversarial attacks for RL
where $\norm{\cdot}$ can be any $p$-norm with $1\leq p\leq\infty$. If the reward function $r_{t}$ and dynamic model $f$ are both differentiable and known to the attacker, (4) can be solved by gradient descent methods. If part of the dynamical system is unknown or not differentiable, solving (4) is difficult or impossible and it is common instead to solve the simplified problem:
@@ 2.3 Lipschitz-bounded deep networks
A simple Lipschitz bound estimation is the layer-wise spectral norm bound $\overline{\gamma}_{s}=\prod_{k=1}^{L}\|W_{k}\|_{2}$. This approach utilizes the fact that common activation functions $\sigma$ are 1-Lipschitz. Thus, we can construct Lipschitz policy networks via scaling factors and 1-Lipschitz linear layers, where three representative examples are given as follows.
@@
A much tighter bound estimation method was recently proposed in [8], which explores both the monotonicity and Lipschitz properties of the activation $\sigma$ by leveraging the integral quadratic constraint (IQC) framework [15]. Direct (i.e. unconstrained) parameterizations based on IQC were proposed in [21] for deep equilibrium networks, in [2] for residual networks, and in [30] for deep MLPs and CNNs. In particular, [30] shows that the IQC-based approach can lead to a 1-Lipschitz nonlinear layer as follows.
@@
    Note that (10) can be transformed back to the MLP form (6). In [30], experimental results on image datasets show that (10) can achieve better performance than 1-Lipschitz linear layers such as AOL [3] and Cayley [28]. It is natural to ask whether similar results hold true for Deep RL.
@@ 3 Experimental Setup
We study two classic RL problems — pendulum swing-up and Atari Pong — to investigate the research questions outlined in Section 1. Our code and training details are available on ${}^{\text{,}}$ [^2] https://github.com/nic-barbara/Lipschitz-RL-Atari  and in Appendix A, respectively.
@@ 4 Results and Discussion
We first study the advantages of Lipschitz-bounded policy networks in terms of robustness to perturbations and adversarial attacks, using pendulum swing-up as an illustrative example (Sec. 4.1). We then extend this study to vision-based feedback control in Pong (Sec. 4.2) and compare the benefits of different Lipschitz-bounded policy architectures. In Figures 4, 5, and Table 1, Lipschitz lower bounds $\underline{\gamma}$ were computed for each policy by performing gradient ascent on the model inputs to maximize the local Lipschitz constant.
@@ 4.2 Comparing architectures — Atari Pong
It is interesting to look deeper into the effect of adversarial attacks on these models. Figure 6 shows just how much of an improvement Lipschitz-bounded policy networks provide in Pong over a standard, unconstrained CNN. The CNN loses the game when subject to very small amounts of random noise and almost imperceptible adversarial attacks. In contrast, the Lipschitz-bounded policy is only beaten with a level of random noise that would even make the game difficult for a human. Moreover, highly-structured $\ell_{2}$-constrained attacks are required to beat the Lipschitz-bounded policy. Looking closely at Figure 6, successful $\ell_{2}$ PGD attacks try to trick the policy into thinking the opponent’s paddle is higher than it actually is while also trying to hide the exact location of the puck. The $\ell_{2}$ attacks also seem to focus on the white boundary walls of the game. It is less clear why the policies should be sensitive to these features, but we hypothesize that the straight lines of the walls may appear similar to a paddle in feature space after passing through convolutional layers. There is no clear structure to the $\ell_{\infty}$ PGD attacks and they remain rather small, since the Lipschitz-bounded policies have a constrained $\ell_{2}$ Lipschitz bound, and the $\ell_{2}$ norm is only a loose upper bound of the $\ell_{\infty}$ norm in high-dimensional spaces. The additional robustness to $\ell_{\infty}$ PGD attacks over CNN policies is a nice bonus.
@@
So far, we have compared Lipschitz-bounded policy networks constructed from Sandwich layers with unconstrained networks to illustrate that smoother policies can improve robustness in Deep RL. It turns out that in addition to tuning its upper bound $\gamma$, the layer architecture we use to bound the Lipschitz constant $\mathrm{Lip}(\kappa)$ of a policy $\kappa$ is extremely important. Figure 7 compares the training curves of all four Lipschitz-bounded policy architectures introduced in Section 2.3 with unconstrained CNN policies, while Table 1 summarizes their nominal and robust performance. It is immediately clear from Figures 7(a) and 7(b) that the two layer architectures known to have conservative bounds on $\mathrm{Lip}(\kappa)$, SN and AOL, perform poorly when $\gamma$ is small. When $\gamma\leq 20$, neither architecture ever produces a winning policy. For larger $\gamma$, the SN and AOL policies can learn to win the game, but the training dynamics are extremely slow in comparison to the CNN (this is a known problem for AOL, see [3, Sec.7]). Interestingly, Table 1 shows that the estimated lower bound $\underline{\gamma}$ on $\mathrm{Lip}(\kappa)$ for these policies is very small when $\gamma$ is small. This suggests that the conservative parameterization of SN and AOL layers restricts their parameter spaces to a small set of very smooth models which does not include high-performing policy networks. In contrast, Figures 7(c) and 7(d) show that the two layer architectures with much tighter bounds on $\mathrm{Lip}(\kappa)$, Cayley and Sandwich, perform quite well even for small choices of $\gamma$. The choice of $\gamma$ still seems to have a strong impact on the training dynamics for these layers — as $\gamma$ increases, so too does the speed at which the policies converge on a winning strategy. This raises interesting questions about the coupling between a policy’s Lipschitz constant and the exploration of its parameter space, which we leave for future work.
@@
Looking closer at Table 1 reveals that simply having a tight Lipschitz bound (i.e., close lower and upper bounds $\underline{\gamma}$ and $\gamma$) is not the only factor contributing to a policy’s performance and robustness. Instead, there appears to be an advantage to using “expressive” policy networks, particularly when high robustness (small $\gamma$) is required. Table 1 indicates that both the Sandwich and Cayley policies exhibit strong performance and robustness for $\gamma\geq 20$. Either layer architecture is therefore a suitable choice for reasonable improvement over existing methods in robust RL like SN. For smaller $\gamma<20$, however, Sandwich policies are far superior, and with $\gamma=10$ they are the most robust of any policy across all three input perturbations while still achieving a strong unperturbed reward of 19.6. This is despite the fact that Cayley policies often have a tighter Lipschitz bound than the Sandwich policies (take $\gamma=10$ as an example). We suggest that this is due to the less conservative parameterization of the Sandwich layers. Each of the SN, AOL, and Cayley policies are composed of linear layers with a spectral norm of approximately 1 (exactly 1 for Cayley). In contrast, Sandwich layers are nonlinear, have no direct restriction on their spectral norm [30, Fig. 4], and instead constrain $\mathrm{Lip}(\kappa)$ via a composition of nonlinear layers which are a complete parameterization of the tightest known bounds on the Lipschitz constant of DNNs [8]. This allows the Sandwich models to converge on policy networks that are both performant and robust even when their parameter space is restricted by a small value of $\gamma$, allowing finer control over the performance-robustness trade-off in Deep RL.
@@ 5 Conclusions
This paper has studied the robustness benefits of Lipschitz-bounded policy networks in Deep RL. We have found that policy networks with small Lipschitz bounds are significantly more robust to perturbations such as disturbances, random noise, and targeted adversarial attacks. Moreover, we have observed that choosing a policy with non-conservative Lipschitz bounds and an expressive, nonlinear layer architecture gives the user finer control over the performance-robustness trade-off than existing methods based on spectral normalization. This raises interesting questions for future study, such as whether Lipschitz-bounded policy networks can augment or alleviate the need for adversarial training, and whether the observed benefits can be transferred to real-world robotic systems.
@@ Acknowledgments
Appendix A Training details
A.1 Pendulum swing-up experiments
The pendulum was modeled in MuJoCo XLA (MJX), a fully-differentiable implementation of the MuJoCo physics simulator [27] written in JAX [5] that allows users to take gradients through the entire closed-loop system. We trained unconstrained MLP and Lipschitz-bounded (Sandwich) policies using the PPO implementation in [9], which leverages the scalability of JAX to massively multi-thread parallel physics simulations on a single workstation GPU. Hyperparameters were tuned by varying each parameter one at a time and choosing the best-performing parameters for an unconstrained MLP. The same hyperparameters were used to train Lipschitz-bounded policies without any further tuning. Our chosen hyperparameters can be found on our GitHub repository [^3] https://github.com/nic-barbara/Lipschitz-RL-MJX/blob/main/scripts/pendulum/train˙2˙train˙models.py. We trained 10 policies for each model architecture and choice of $\gamma$, each with a different random seed for model initialization. Results showing robust performance to sample delays and adversarial attacks in Figure 4 were averaged over the 10 policies and 1024 pendulum environments starting from random initial states. MLP networks were composed of 4 linear layers of 32 hidden nodes. Lipschitz-bounded policies were composed of 4 Sandwich layers of 21 hidden nodes to ensure the two model architectures had a similar number of trainable parameters. We chose tanh activations for all policies in accordance with [1].
A.2 Atari Pong experiments
We trained unconstrained CNN models and the Lipschitz-bounded policy architectures (SN, AOL, Sandwich, Cayley) across 4 random model initialisations using the PPO implementation in [10] and the ALE/Pong-v5 environment from [31]. We used the default hyperparameters chosen for CNN policies in [10] for all policy architectures. We could not use the reward gradient to directly optimize (4) as the Pong environment is not differentiable, hence we implemented attacks with the PGD method [7]. Robust performance to noise and adversarial attacks in Figure 5 and Table 1 were averaged over the 4 policies and 20 games of Pong, each with a different random seed. Further details on the network architecture for each layer type can be found on our GitHub repository [^4] https://github.com/nic-barbara/Lipschitz-RL-Atari/blob/main/liprl/atari˙agent.py.

### v3

@@ Abstract
This paper presents a study of robust policy networks in deep reinforcement learning. We investigate the benefits of policy parameterizations that naturally satisfy constraints on their Lipschitz bound, analyzing their empirical performance and robustness on two representative problems: pendulum swing-up and Atari Pong. We illustrate that policy networks with smaller Lipschitz bounds are more robust to disturbances, random noise, and targeted adversarial attacks than unconstrained policies composed of vanilla multi-layer perceptrons or convolutional neural networks. However, the structure of the Lipschitz layer is important. We find that the widely-used method of spectral normalization is too conservative and severely impacts clean performance, whereas more expressive Lipschitz layers such as the recently-proposed Sandwich layer can achieve improved robustness without sacrificing clean performance.
@@
reinforcement learning robustness Lipschitz networks
@@ 1 Introduction
Deep reinforcement learning (deep RL) has been the driving force behind many recent successes in learning-based control, including in discrete game-like problems [15], robotic manipulation [10], and locomotion [22]. However, the applicability of deep RL to performance- and safety-critical systems is currently limited by questions of its robustness [8]. It is well known that neural networks can be highly sensitive to small input perturbations [25], making policy networks learned via deep RL potentially unrobust to disturbances, noise, and targeted adversarial attacks. Despite sharing similar sensitivity issues to neural classifiers, for which many robust neural networks have recently been developed [20, 28, 30], the use of robust policy architectures in deep RL has not been widely studied.
@@
The true $\ell_{2}$ Lipschitz constant, denoted $\mathrm{Lip}(f)$, is the smallest $\gamma$ satisfying (1). Lipschitz-bounded networks are “smoother” and less sensitive to input perturbations than unconstrained networks since small variations in their inputs only induce small variations in their outputs. Despite the wealth of recent work on constructing Lipschitz-bounded deep networks [19, 28, 30], to the best of our knowledge, the only method tested in deep RL is spectral normalization [3, 26] which is known to give conservative bounds on the true Lipschitz constant.
@@ 2.1 Deep Reinforcement Learning
$\max_{\theta}\ R:=\mathop{\mathbb{E}}_{x_{0},w}\left[\sum_{t=0}^{\infty}\rho^{t}r_{t}(x_{t},u_{t})\right]\,\text{s.t.}\,\ x_{t+1}=f(x_{t},u_{t},w_{t}),\ u_{t}=\kappa(x_{t};\theta),$ (3)
@@ 2.2 Adversarial Attacks for RL
where $\norm{\cdot}$ can be any $p$-norm with $1\leq p\leq\infty$. If the reward function $r_{t}$ and dynamic model $f$ are both differentiable and known to the attacker, (4) can be solved by gradient descent methods. If part of the dynamical system is unknown or not differentiable, solving (4) is difficult or impossible and it is common instead to solve the simplified problem
@@ 2.3 Lipschitz-Bounded Deep Networks
A simple Lipschitz bound estimation is the layer-wise spectral norm bound $\overline{\gamma}_{s}=\prod_{k=1}^{L}\|W_{k}\|_{2}$. This approach uses the fact that common activation functions $\sigma$ are 1-Lipschitz. Thus, we can construct Lipschitz policy networks via scaling factors and 1-Lipschitz linear layers, where three representative examples are given as follows.
@@
A tighter and more expressive bound estimation method was recently proposed in [6], which explores both the monotonicity and Lipschitz properties of the activation $\sigma$ by leveraging the integral quadratic constraint (IQC) framework [13]. Direct (i.e. unconstrained) parameterizations based on IQC were proposed in [21] for deep equilibrium networks, in [2] for residual networks, and in [30] for deep MLPs and CNNs. In particular, [30] shows that the IQC-based approach can lead to a 1-Lipschitz nonlinear layer as follows.
@@
    Note that (10) can be transformed back to the MLP form (6).
The Sandwich layer (9) is an “expressive” layer architecture in that it contains all 1-Lipschitz linear layers as a special case, and it allows the spectral norm bound of each layer (and their product) to be greater than one [30]. Experimental results in [30] on image datasets show that (10) can achieve better performance than 1-Lipschitz linear layers such as AOL and Cayley. It is natural to ask whether similar results hold true for deep RL.
@@ 3 Experimental Setup
We study two classic RL problems — pendulum swing-up and Atari Pong — to investigate the research questions outlined in Section 1. Our code and training details are available on ${}^{\text{,}}$ [^2] https://github.com/nic-barbara/Lipschitz-RL-Atari  and in Appendix 0.A, respectively.
@@ 4 Results and Discussion
We first study the advantages of Lipschitz-bounded policy networks in terms of robustness to perturbations and adversarial attacks, using pendulum swing-up as an illustrative example (Sec. 4.1). We extend this study to vision-based feedback control in Pong (Sec. 4.2) and compare the benefits of different Lipschitz-bounded policy architectures. In Figures 4, 5, and Table 1, Lipschitz lower bounds $\underline{\gamma}$ were computed for each policy by performing gradient ascent on (8).
@@ 4.2 Comparing Architectures — Atari Pong
It is interesting to look deeper into the effect of adversarial attacks on these models. Figure 6 shows just how much of an improvement Lipschitz-bounded policy networks provide in Pong over a standard, unconstrained CNN. The CNN loses the game when subject to very small amounts of random noise and almost imperceptible adversarial attacks. In contrast, the Lipschitz-bounded policy is only beaten with a level of random noise that would even make the game difficult for a human. Moreover, highly-structured $\ell_{2}$-constrained attacks are required to beat the Lipschitz-bounded policy. Looking closely at Figure 6, successful $\ell_{2}$ PGD attacks try to trick the policy into thinking the opponent’s paddle is shifted from where it actually is while also trying to hide the exact location of the puck. The $\ell_{2}$ attacks also seem to focus on the white boundary walls of the game. It is less clear why the policies should be sensitive to these features, but we hypothesize that the straight lines of the walls may appear similar to a paddle in feature space after passing through convolutional layers. There is no clear structure to the $\ell_{\infty}$ PGD attacks and they remain rather small, since the Lipschitz-bounded policies have a constrained $\ell_{2}$ Lipschitz bound, and the $\ell_{2}$ norm is only a loose upper bound of the $\ell_{\infty}$ norm in high-dimensional spaces. The additional robustness to $\ell_{\infty}$ PGD attacks over CNN policies is a nice bonus.
@@
So far, we have only investigated Lipschitz-bounded policy networks constructed from Sandwich layers to illustrate that smoother policies can improve robustness in deep RL. It turns out that in addition to choosing its upper bound $\gamma$, the layer architecture we use to bound the Lipschitz constant $\mathrm{Lip}(\kappa)$ of a policy $\kappa$ is extremely important. Figure 7 compares the training curves of all four Lipschitz-bounded policy architectures from Section 2.3 with unconstrained CNN policies, while Table 1 summarizes their nominal and robust performance. It is immediately clear from Figures 7(a) and 7(b) that the two layer architectures known to have conservative bounds, SN and AOL, perform poorly when $\gamma$ is small — e.g., when $\gamma\leq 20$, neither architecture ever produces a winning policy. For larger $\gamma$, the SN and AOL policies learn to win the game, but their training dynamics are extremely slow in comparison to the CNN (this is a known problem for AOL, see [20, Sec.7]). Moreover, Table 1 shows that the estimated lower bound $\underline{\gamma}$ on $\mathrm{Lip}(\kappa)$ for these policies is very small when $\gamma$ is small. This suggests that the conservative parameterization of SN and AOL layers restricts their parameter spaces to a small set of very smooth models which does not include high-performing policies. In contrast, Figures 7(c) and 7(d) show that the two layer architectures with much tighter bounds on $\mathrm{Lip}(\kappa)$, Cayley and Sandwich, perform quite well even for small choices of $\gamma$. The choice of $\gamma$ still seems to have a strong impact on the training dynamics for these layers — as $\gamma$ increases, so too does the speed at which the policies converge on a winning strategy. This raises interesting questions about the coupling between a policy’s Lipschitz constant and the exploration of its parameter space, which we leave for future work.
@@
Looking closer at Table 1 reveals that simply having a tight Lipschitz bound (i.e., close lower and upper bounds $\underline{\gamma}$ and $\gamma$) is not the only factor contributing to a policy’s performance and robustness. Instead, there appears to be an advantage to using “expressive” policy networks, particularly when high robustness (small $\gamma$) is required. Table 1 indicates that both the Sandwich and Cayley policies exhibit strong performance and robustness for $\gamma\geq 20$. Either layer architecture is therefore a suitable choice for reasonable improvement over existing methods in robust RL like SN. For smaller $\gamma<20$, however, Sandwich policies are far superior, and with $\gamma=10$ they are the most robust of any policy across all three input perturbations while still achieving a strong unperturbed reward of 19.6. This is despite the fact that Cayley policies often have a tighter Lipschitz bound than the Sandwich policies (take $\gamma=10$ as an example). We suggest that this is due to the less conservative parameterization of the Sandwich layers. Each of the SN, AOL, and Cayley policies are composed of linear layers with a spectral norm of approximately 1 (exactly 1 for Cayley). In contrast, Sandwich layers have no direct restriction on their spectral norm [30, Fig. 4], and instead constrain $\mathrm{Lip}(\kappa)$ via a composition of nonlinear layers which are a complete parameterization of all networks satisfying the tightest known bounds on the Lipschitz constant of DNNs [6], of which 1-Lipschitz linear layers are a special case. This expressivity allows the Sandwich models to converge on policy networks that are both performant and robust even when their parameter space is restricted by a small value of $\gamma$, allowing finer control over the performance-robustness trade-off in deep RL.
@@ 5 Conclusions
This paper has studied the robustness benefits of Lipschitz-bounded policy networks in deep RL. We have found that policy networks with small Lipschitz bounds are significantly more robust to perturbations such as disturbances, random noise, and targeted adversarial attacks. Moreover, we have observed that choosing a policy with non-conservative Lipschitz bounds and an expressive layer architecture gives the user finer control over the performance-robustness trade-off than existing methods based on spectral normalization. This raises interesting questions for future study, such as whether Lipschitz-bounded policy networks can augment or alleviate the need for adversarial training, and whether the observed benefits can be transferred to real-world robotic systems.
{credits}
@@ Acknowledgments
Appendix 0.A Training details
0.A.1 Pendulum swing-up experiments
The pendulum was modeled in MuJoCo XLA (MJX), a fully-differentiable implementation of the MuJoCo physics simulator [27] written in JAX [4] that allows users to take gradients through the entire closed-loop system. We trained unconstrained MLP and Lipschitz-bounded (Sandwich) policies using the PPO implementation in [7], which leverages the scalability of JAX to massively multi-thread parallel physics simulations on a single workstation GPU. Hyperparameters were tuned by varying each parameter one at a time and choosing the best-performing parameters for an unconstrained MLP. The same hyperparameters were used to train Lipschitz-bounded policies without any further tuning. Our chosen hyperparameters can be found on our GitHub repository [^3] https://github.com/nic-barbara/Lipschitz-RL-MJX/blob/main/scripts/pendulum/train_2_train_models.py. We trained 10 policies for each model architecture and choice of $\gamma$, each with a different random seed for model initialization. Results showing robust performance to sample delays and adversarial attacks in Figure 4 were averaged over the 10 policies and 1024 pendulum environments starting from random initial states. MLP networks were composed of 4 linear layers of 32 hidden nodes. Lipschitz-bounded policies were composed of 4 Sandwich layers of 21 hidden nodes to ensure the two model architectures had a similar number of trainable parameters. We chose tanh activations for all policies in accordance with [1].
0.A.2 Atari Pong experiments
We trained unconstrained CNN models and the Lipschitz-bounded policy architectures (SN, AOL, Sandwich, Cayley) across 4 random model initialisations using the PPO implementation in [9] and the ALE/Pong-v5 environment from [31]. We used the default hyperparameters chosen for CNN policies in [9] for all policy architectures. We could not use the reward gradient to directly optimize (4) as the Pong environment is not differentiable, hence we implemented attacks with the PGD method [5]. Robust performance to noise and adversarial attacks in Figure 5 and Table 1 were averaged over the 4 policies and 20 games of Pong, each with a different random seed. Further details on the network architecture for each layer type can be found on our GitHub repository [^4] https://github.com/nic-barbara/Lipschitz-RL-Atari/blob/main/liprl/atari_agent.py.